Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Comber Open 2001

Distribution

Dr. Trevor West DUCAC Chairman
Emma Stokes DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue Comber Rifle Club, Comber
Date August 4-5, 2001
Discipline 50m Prone, 100yd Prone

DURC Attendees

Steven Watterson Mark Dennehy
John Keeney  Peter Byrne

Ray Reilly Scores 100 Yards

Steven Watterson 384
Ray Reilly 380
John Keeney 383
Peter Byrne 384
Mark Dennehy 342

50 Metres

Steven Watterson 577
Ray Reilly 562
John Keeney 570
Peter Byrne 569
Mark Dennehy 536

Teams Competition DURC Angles

Mark Dennehy 567
John Keeney 583
Peter Byrne 578
Ray Reilly 579
Total 2307

Third Place

North&South

Steven Watterson 581
Total 2332

First Place

Prizes Won

Steven Watterson First place, teams competition

Eight place overall

John Keeney 3rd, Class B, 50m
Peter Byrne 2nd, Class B, 100yds

Report

The Comber Rifle Club is renowned throughout Ireland and the UK as being the toughest range in either island to shoot on. Situated so close to the shore, tidal winds are an enormous challange, and the topography of the range means that wind conditions can change from moment to moment and one firing point can have a totally different wind direction from another only a few feet away. As a result, the 50m Open held in Comber each year is perhaps the most challenging Ireland has to offer and it is little surprise that shooters from Comber such as Olympian Gary Duff, are heavily featured amongst Ireland’s best.

Comber Rifle Club

DURC competeted in this year’s Comber Open with a squad of five shooters. Logistics proved to be a stumbling block, with far too much equipment for one car. Olympian Rhona Barry kindly helped out by ferrying equipment and one of our shooters. The conditions were typical for Comber - changing winds, changing light levels and occasional torrents of rain.

All of the DURC shooters acquitted themselves well, except perhaps for myself, though I had more than my fair share of problems - the sights for my rifle were accidentally not packed so I had to jury-rig a sights setup, and to add insult to injury I was bitten by some insect and had an allergic reaction to the bite.

Steven managed to display a total lack of impediment however, and made it to the finals round, though he didn’t place for a medal.

John Keeney and Ray Reilly

Steven Watterson reflecting between cards during the finals

The finals; and John and Steven relaxing after a long day of shooting

Misé le meas,

Mark Dennehy, DURC Secretary

4